# TC-GS: A Faster Gaussian Splatting Module Utilizing Tensor Cores

### v1

@@ 3.2 Runtime Analysis
Figure 2: a) Three types of fragments on a pixel. b) If the Gaussian only covers a small portion of tile, a large amount culled fragments are generated.
@@
Table 2: The alpha computation and culling steps are run on both blended and culled fragments, while the blending step is only performed on blended fragments. There is no computation on skipped fragments.
@@
$\mathcal{C}=k_{\text{blend}}f_{\text{blend}}+(k_{\text{cull}}+k_{\alpha})(f_{\text{blend}}+f_{\text{cull}}).$ (7)
@@
Figure 3: Analysis of 3DGS rendering bottlenecks: The left side shows the time distribution of preprocessing, sorting, and alpha-blending for 3DGS, AdR-Gaussian, and Speedy-Splat, with alpha-blending dominating. The right side details alpha computation, culling, and blending, identifying culled fragments as the primary bottleneck due to redundant alpha computations, while skipped fragments incur no cost.
4 Method
Figure 4: Design of frag2Mat: the alpha computation is reformulated as matrix multiplication to fully leverage Tensor Cores for accelerating alpha calculation.
This section introduces how TC-GS utilizes the Tensor Cores to accelerate the alpha-computation and culling process, which are the bottlenecks of rendering. The fundamental components of TC-GS include frag2Mat for matrix transformation and G2L for coordinate transformation. The detailed design is shown in Figure 4.
4.1 Batched Exponents Computation with Tensor Cores: Fitting the MMA Operation
@@
We propose frag2Mat, a computational restructuring that transforms per-pixel-per-Gaussian exponent calculations into unified matrix operations. By expressing the exponent as a linear combination of pixel coordinates, Gaussian parameters and their quadratic terms, we decouple multivariate dependencies into two independent components: a pixel vector encoding coordinate polynomials for its pixel, and a Gaussian vector composed of parameters from its Gaussian. The final exponent values are batch-calculated through standard matrix multiplication of two matrices formed by stacking pixel vectors and Gaussian vectors respectively.
Firstly, TC-GS converts Equation 4 into $\alpha_{i}^{(j)}=\exp\left({\beta_{i}^{(j)}}\right)$, where
$\beta_{i}^{(j)}=\ln(o^{(j)})-\frac{1}{2}(\mu^{\prime(j)}-p_{i})^{T}(\Sigma^{\prime(j)})^{-1}(\mu^{\prime(j)}-p_{i}).$ (8)
The core idea of TC-GS is decoupling the computation of exponents $\beta_{i}^{(j)}$ into two separated terms: one that depends only on the pixel index and another that depends only on the Gaussian. Notably, we observe that the exponent can be expressed as a quadratic function of the pixel $p$. Based on this observation, TC-GS transforms Equation 8 into the following standard from:
$\beta_{i}^{(j)}=v_{0}^{(j)}+v_{1}^{(j)}p_{x_{i}}+v_{2}^{(j)}p_{y_{i}}+v_{3}^{(j)}p_{x_{i}}^{2}+v_{4}^{(j)}p_{x_{i}}p_{y_{i}}+v_{5}^{(j)}p_{y_{i}}^{2},$ (9)
@@
$v_{0}^{(j)}$ $=\ln(o^{(j)})-\frac{1}{2}(\mu^{\prime(j)})^{T}(\Sigma^{\prime(j)})^{-1}\mu^{\prime(j)},$ (10)
$v_{1}^{(j)}$ $=a^{(j)}\mu_{x}^{(j)}+b^{(j)}\mu_{y}^{(j)},\quad v_{2}^{(j)}=b^{(j)}\mu_{x}^{(j)}+c^{(j)}\mu_{y}^{(j)},$
$v_{3}^{(j)}$ $=-\dfrac{1}{2}a^{(j)},\quad v_{4}^{(j)}=-b^{(j)},\quad v_{5}^{(j)}=-\dfrac{1}{2}c^{(j)}$
with the notation that $\Sigma^{\prime-1}=\begin{bmatrix}a&b\\ b&c\end{bmatrix}$. The Equation 9 is the expanded dot product of two vectors:
$\beta_{i}^{(j)}=\mathbf{u}_{i}^{T}\mathbf{v}^{(j)},$ (11)
@@
$\mathbf{u}_{i}=\begin{bmatrix}1&p_{x_{i}}&p_{y_{i}}&p_{x_{i}}^{2}&p_{x_{i}}p_{y_{i}}&p_{y_{i}}^{2}\end{bmatrix}^{T}$ (12)
@@
$\mathbf{v}^{(j)}=\begin{bmatrix}v_{0}^{(j)}&v_{1}^{(j)}&v_{2}^{(j)}&v_{3}^{(j)}&v_{4}^{(j)}&v_{5}^{(j)}\end{bmatrix}.$ (13)
Here $\mathbf{u}_{i}$, the pixel vector solely dependent on pixel properties, which will be computed only once for each pixel. Similarly, each Gaussian vector $\mathbf{v}^{(j)}$ can be only computed once. Therefore, the expression of the exponent is decoupled into the dot product of two individual vectors. To align the computation to the Tensor Cores, TC-GS stacks the vectors into two independent matrices:
$\mathbf{U}=\begin{bmatrix}\mathbf{u}_{1}&\mathbf{u}_{2}&\cdots&\mathbf{u}_{m}\end{bmatrix}\in\mathbb{R}^{6\times m},$ (14)
$\mathbf{V}=\begin{bmatrix}\mathbf{v}^{(1)}&\mathbf{v}^{(2)}&\cdots&\mathbf{v}^{(n)}\end{bmatrix}\in\mathbb{R}^{6\times n},$ (15)
@@
$\mathbf{B}=\begin{bmatrix}\beta_{1}^{(1)}&\cdots&\beta_{1}^{(n)}\\ \vdots&\ddots&\vdots\\ \beta_{m}^{(1)}&\cdots&\beta_{m}^{(n)}\\ \end{bmatrix}=\begin{bmatrix}\mathbf{u}_{1}^{T}\mathbf{v}^{(1)}&\cdots&\mathbf{u}_{1}^{T}\mathbf{v}^{(n)}\\ \vdots&\ddots&\vdots\\ \mathbf{u}_{m}^{T}\mathbf{v}^{(1)}&\cdots&\mathbf{u}_{m}^{T}\mathbf{v}^{(n)}\\ \end{bmatrix}=\mathbf{U}^{T}\mathbf{V}.$ (16)
@@
4.2 Transform Global to Local Coordinates: Fitting the Half Precision
Although we adopted Tensor Cores for alpha computation and culling. The challenges remain as Tensor Cores do not support common single-precision floating-point operations. Instead, they only take FP16 or TF32 as the input of the MMA operation, resulting in a higher machine epsilon of $\epsilon_{H}=9.77\times 10^{-4}$ compared to FP32’s $\epsilon_{F}=1.19\times 10^{-7}$. The rounding error of a floating number $r$ is $|r|\times\epsilon_{H}$. Therefore, the error of computation on Tensor Cores is more sensitive at the absolute value of the input data.
We notice that the range of the pixel $p$ is $[0,w]\times[0,h]$, where $w,h$ are the width and height of the screen space, respectively. When loading the pixel vector $\mathbf{u}$ in Equation 12 into Tensor Cores, its absolute values of quadratic terms $p_{x}^{2}$, $p_{x}p_{y}$ and $p_{y}^{2}$ can exceed a million, contributing catastrophic rounding errors to $\beta$, whose relative error is more than $35000\%$ at the resolution of 720p as proved on Appendix B.1. Moreover, overflow will occur if the FP16 range $[-6.55\times 10^{5},6.55\times 10^{5}]$ is used.
Figure 5: The quadratic term of pixel coordinates significantly contributes the rounding error, resulting in blurry images. Local coordinates constrain the value of $\Delta p$ into $[-8,8]^{2}$, which can reduce the upper bound of the rounding error.
@@ 4.2 Transform Global to Local Coordinates: Fitting the Half Precision
$\Delta\mu^{\prime}=\mu^{\prime}-p(\mathcal{T}).$ (18)
@@
$\Delta\mu^{\prime}-\Delta p=\mu^{\prime}-p.$ (19)
Therefore, frag2Mat can be also performed on local coordinates. As the proof in Appendix B goes, G2L reduces the quadratic rounding error $O(h^{2}+hw+w^{2})$ into linear $O(h+w)$, and achieve better rendering quality in FP16 as Section 5.3 demonstrates.
4.3 Implementation Details
TC-GS also incorporates several common optimization techniques to further enhance overall performance.
First, TC-GS adopts an early-pruning strategy: it identifies invisible Gaussians that can be culled before the alpha-computation stage, greatly reducing the number of exponential (exp) instructions executed and improving throughput.
We begin by expressing the exponentiation in logarithmic form, allowing us to precompute each Gaussian’s opacity during the preprocessing stage. This lets us filter out non-contributing Gaussians in advance, reducing the amount of exp instructions issued by alpha computation. Since prior experiments show that culled fragments account for a large portion of the workload, this approach yields a substantial speed-up for the entire pipeline.
Additionally, TC-GS precisely arranges the shared-memory layout to minimize bank conflicts during warp-level shared-memory reads and writes. This maximizes the operator’s effective bandwidth and fully unleashes Tensor Core performance.
5 Experiment
5.1 Experiment Setup
Table 3: Comparison With/Without TC-GS Module Across Datasets
@@ 5.1 Experiment Setup
AdR-Gaussian(+TC-GS) 545.154(1.955$\times$) 23.632 0.851 0.169 612.063(1.89$\times$) 29.815 0.906 0.236 467.310(1.78$\times$) 26.519 0.785 0.25
@@
To prove that our method is universal and efficient, we conduct experiments on the same datasets as those used in 3DGS[13]. Specifically, the datasets contain all scenes from Mip-NeRF360[22], two outdoor scenes (Truck & Train) from Tanks & Temple[23], two indoor scenes (Drjohnson & playroom) in Deep-Blending[24]. For each scene, CUDA events are inserted at the start and end of the forward rendering procedure to measure FPS values, and the reported results represent the average metrics of the test datasets.All experiments are conducted on an NVIDIA A800 GPU (80GB SXM) equipped with 432 Tensor Cores, achieving a peak 624 TFLOPS FP16 performance.
5.2 Comparisons
@@
Across all datasets, incorporating the TC-GS module roughly doubles frame throughput except for FlashGS. The state-of-the-art performance is attained by FlashGS integrated with TC-GS, which achieves a 3.3–5.6$\times$ speedup compared to the original 3DGS, while delivering rendering speeds of 479.3–735.8 FPS.
By comparing the performance of alpha-blending, as shown in Figure 4, the rendering pipeline gains a 2-4$\times$ performance improvement. Combined with the previous analysis, this further emphasizes the importance of optimizing the rendering pipeline.
While AdR-Gaussian and Speedy-Splat already employ advanced optimizations to reduce redundant computation, TC-GS further pushed their performance boundaries, delivering significant speedup of 1.87$\times$ and 1.84$\times$ respectively. As for FlashGS, while its existing pipeline rendering optimizations yield a more modest 1.38$\times$ average acceleration with TC-GS integration, we emphasize that the pipeline optimizations and TC-GS’s approach operate orthogonally.
@@
Table 4: Comparison of alpha-blending time across methods and datasets.
@@
3DGS drjohnson 1.348(3.49$\times$) 4.705
train 1.217(3.57$\times$) 4.348
flowers 0.880(3.68$\times$) 3.236
FlashGS drjohnson 0.540(2.18$\times$) 1.179
train 0.609(2.03$\times$) 1.234
flowers 0.625(2.38$\times$) 1.487
AdR-Gaussian drjohnson 0.406(4.76$\times$) 1.931
train 0.518(4.23$\times$) 2.191
flowers 0.469(4.39$\times$) 2.061
Speedy-Splat drjohnson 0.510(3.86$\times$) 1.969
train 0.672(3.46$\times$) 2.326
flowers 0.592(3.52$\times$) 2.083
5.3 Ablation Study
To clearly illustrate the impact of G2L on rendering quality and performance, we ran the following comparisons shown in Table 5 and the rendering images are shown in Fig.6. The ablation experiment shows that all Tensor Core variants boosted FPS. However, direct TF32 computation already introduced noticeable image-quality degradation, and FP16 alone dropped PSNR to just 8 dB. In contrast, combining FP16 Tensor Core optimization with Local Coordinates method preserved nearly same image quality while delivering the highest FPS. This demonstrates that our G2L can simultaneously maximize performance and maintain rendering fidelity.
Table 5: Ablation study on applying TC-GS on original 3DGS.
frag2Mat G2L Precison drjohnson truck
@@ 5.3 Ablation Study
Figure 6: Ablation study on G2L when applying TC-GS on original 3DGS.
5.4 Discussion
5.4.1 Limitations

### v2

@@ 3.2. Observation and Motivation
Figure 2. a) Three types of fragments on a pixel. b) If the Gaussian only covers a small portion of tile, a large amount culled fragments are generated.
@@
Table 2. The alpha computation and culling steps are run on both blended and culled fragments, while the blending step is only performed on blended fragments. There is no computation on skipped fragments.
@@
(7) $\mathcal{C}=k_{\text{blend}}f_{\text{blend}}+(k_{\text{cull}}+k_{\alpha})(f_{\text{blend}}+f_{\text{cull}}).$
@@
Figure 3. Analysis of 3DGS rendering bottlenecks: The left side shows the time distribution of preprocessing, sorting, and alpha-blending for 3DGS, AdR-Gaussian, and Speedy-Splat, with alpha-blending dominating. The right side details alpha computation, culling, and blending, identifying culled fragments as the primary bottleneck due to redundant alpha computations, while skipped fragments incur no cost.
4. Method
Figure 4. Design of Frag2Mat: the alpha computation is reformulated as matrix multiplication to fully leverage Tensor Cores for accelerating alpha calculation.
This section introduces how TC-GS utilizes the Tensor Cores to accelerate the alpha-computation and culling process, which are the bottlenecks of rendering. The fundamental components of TC-GS include EarlyCull for culling optimization, Frag2Mat for matrix transformation and G2L for coordinate transformation. The detailed design is shown in Figure 4.
4.1. EarlyCull: Reducing Exponential Computation
First, TC-GS adopts EarlyCull, an early-pruning strategy: it identifies invisible Gaussians that can be culled before the alpha-computation stage, greatly reducing the number of exponential instructions exp executed and improving throughput.
We begin by expressing Equation (4) as the following form:
(8) $\alpha_{i,j}=\mathrm{e}^{\beta_{i,j}},$
where the exponent $\beta_{i,j}$ satisfies
(9) $\beta_{i,j}=\ln(o_{j})-\frac{1}{2}(\mu^{\prime}_{j}-p_{i})^{T}(\Sigma^{\prime}_{j})^{-1}(\mu^{\prime}_{j}-p_{i}).$
Then the culling condition, Equation (5), can be converted into
(10) $\beta_{i,j}<-\ln(255).$
EarlyCull allows us to filter out non-contributing Gaussians in advance, reducing the amount of expensive exp instructions issued by alpha computation. Since prior experiments show that culled fragments account for a large portion of the workload, this approach yields a substantial speed-up for the entire pipeline.
4.2. Frag2Mat: Fitting the MMA Operation
@@
We propose Frag2Mat, a computational restructuring that transforms per-pixel-per-Gaussian exponent calculations into unified matrix operations. By expressing the exponent as a linear combination of pixel coordinates, Gaussian parameters and their quadratic terms, we decouple multivariate dependencies into two independent components: a pixel vector encoding coordinate polynomials for its pixel, and a Gaussian vector composed of parameters from its Gaussian. The final exponent values are batch-calculated through standard matrix multiplication of two matrices formed by stacking pixel vectors and Gaussian vectors respectively. LC The core idea of TC-GS is decoupling the computation of exponents $\beta_{i,j}$ into two separated terms: one that depends only on the pixel index and another that depends only on the Gaussian. Notably, we observe that the exponent can be expressed as a quadratic function of the pixel $p$. Based on this observation, TC-GS transforms Equation (9 into the following standard from:
(11) $\beta=v_{0}+v_{1}p_{x}+v_{2}p_{y}+v_{3}p_{x}^{2}+v_{4}p_{x}p_{y}+v_{5}p_{y}^{2},$
@@
(12) $v_{0}$ $=\ln(o)-\frac{1}{2}\mu^{\prime T}\Sigma^{\prime-1}\mu^{\prime},$
$v_{1}$ $=\sigma_{11}\mu^{\prime}_{x}+\sigma_{12}\mu^{\prime}_{y},\quad v_{2}=\sigma_{12}\mu^{\prime}_{x}+\sigma_{22}\mu^{\prime}_{y},$
$v_{3}$ $=-\dfrac{1}{2}\sigma_{11},\quad v_{4}=-\sigma_{12},\quad v_{5}=-\dfrac{1}{2}\sigma_{22}$
with the notation that symmetric $\Sigma^{\prime-1}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}$ with $\sigma_{12}=\sigma_{21}$.
The Equation (11) is the expanded dot product of two vectors:
(13) $\beta_{i,j}=\mathbf{u}_{i}^{T}\mathbf{v}_{j},$
@@
(14) $\mathbf{u}=\begin{bmatrix}1&p_{x}&p_{y}&p_{x}^{2}&p_{x}p_{y}&p_{y}^{2}\end{bmatrix}^{T}$
@@
(15) $\mathbf{v}=\begin{bmatrix}v_{0}&v_{1}&v_{2}&v_{3}&v_{4}&v_{5}\end{bmatrix}^{T}.$
Here $\mathbf{u}$, the pixel vector solely dependent on pixel properties, which will be computed only once for each pixel. Similarly, each Gaussian vector $\mathbf{v}$ can be only computed once. Therefore, the expression of the exponent is decoupled into the dot product of two individual vectors. To align the computation to the Tensor Cores, TC-GS stacks the vectors into two independent matrices:
(16) $\mathbf{U}=\begin{bmatrix}\mathbf{u}_{1}&\mathbf{u}_{2}&\cdots&\mathbf{u}_{m}\end{bmatrix}\in\mathbb{R}^{6\times m},$
(17) $\mathbf{V}=\begin{bmatrix}\mathbf{v}_{1}&\mathbf{v}_{2}&\cdots&\mathbf{v}_{j}\end{bmatrix}\in\mathbb{R}^{6\times n},$
@@
(18) $\mathbf{B}=\begin{bmatrix}\beta_{1,1}&\cdots&\beta_{1,n}\\ \vdots&\ddots&\vdots\\ \beta_{m,1}&\cdots&\beta_{m,n}\\ \end{bmatrix}=\begin{bmatrix}\mathbf{u}_{1}^{T}\mathbf{v}_{1}&\cdots&\mathbf{u}_{1}^{T}\mathbf{v}_{n}\\ \vdots&\ddots&\vdots\\ \mathbf{u}_{m}^{T}\mathbf{v}_{1}&\cdots&\mathbf{u}_{m}^{T}\mathbf{v}_{n}\\ \end{bmatrix}=\mathbf{U}^{T}\mathbf{V}.$
@@
4.3. G2L: Fitting the Half Precision
Although we adopted Tensor Cores for alpha computation and culling. The challenges remain as Tensor Cores do not support common single-precision floating-point operations. Instead, they only take FP16 or TF32 as the input of the MMA operation, resulting in a higher machine epsilon of $\varepsilon_{H}=9.77\times 10^{-4}$ compared to FP32’s $\varepsilon_{F}=1.19\times 10^{-7}$. The rounding error of a floating number $r$ is $|r|\varepsilon_{H}$. Therefore, the error of computation on Tensor Cores is more sensitive at the absolute value of the input data.
We notice that the range of the pixel $p$ is $0,w]\times[0,h]$, where $w,h$ are the width and height of the screen space, respectively. When loading the pixel vector $\mathbf{u}$ in Equation (14) into Tensor Cores, its absolute values of quadratic terms $p_{x}^{2}$, $p_{x}p_{y}$ and $p_{y}^{2}$ can exceed a million, contributing catastrophic rounding errors to $\beta$. Moreover, overflow will occur if the FP16 range $6.55\times 10^{5},6.55\times 10^{5}]$ is used.
Figure 5. The quadratic term of pixel coordinates significantly contributes the rounding error, resulting in blurry images. Local coordinates constrain the value of $\Delta p$ into $8,8]^{2}$, which can reduce the upper bound of the rounding error.
@@ 4.3. G2L: Fitting the Half Precision
(20) $\Delta\mu^{\prime}=\mu^{\prime}-p(\mathcal{T}).$
@@
(21) $\Delta\mu^{\prime}-\Delta p=\mu^{\prime}-p.$
Therefore, Frag2Mat can be also performed on local coordinates.
As the proof in the appendices goes, G2L reduces the quadratic rounding error $O(h^{2}+hw+w^{2})\varepsilon_{H}$ into linear $O(h+w)\varepsilon_{H}$, and achieve better rendering quality in FP16 as Section 5.3 demonstrates.
5. Experiment
5.1. Experiment Setup
Table 3. Comparison With/Without TC-GS Module Across Datasets.
@@ 5.1. Experiment Setup
[…]
Implementation
Current computation patterns which Tensor Cores support do not include length-6 MMA operations. A simple way to align the required length-8 MMA operation is zero padding. In our implementation, we replace $v_{0}$ in $\mathbf{v}$ with $\frac{v_{0}}{3}$ and duplicate it 3 times. The constant term $1$ in $\mathbf{u}$ are duplicated respectively. This approach can reduce the rounding error by reducing the absolute value.
@@
To prove that our method is universal and efficient, we conduct experiments on the same datasets as those used in 3DGS(Kerbl et al., 2023). Specifically, the datasets contain all scenes from Mip-NeRF360(Barron et al., 2022), two outdoor scenes (Truck & Train) from Tanks & Temple(Knapitsch et al., 2017), two indoor scenes (Drjohnson & playroom) in Deep-Blending(Hedman et al., 2018). For each scene, CUDA events are inserted at the start and end of the forward rendering procedure to measure FPS values, and the reported results represent the average metrics of the test datasets.All experiments are conducted on an NVIDIA A800 GPU (80GB SXM) equipped with 432 Tensor Cores, achieving a peak 624 TFLOPS FP16 performance.
5.2. Comparisons
@@
[…]
@@
Table 4. Comparison of alpha-blending time across methods and datasets.
@@
[…]
5.3. Ablation Study
To determine the source of the acceleration, we conducted an ablation study on EarlyCull and Frag2Mat with G2L across various renderers and datasets. As Table 5 shows, both methods can accelerate the rendering speed. EarlyCull alone achieves $1.06-1.51\times$ speedup, but it accelerates slightly when Frag2Mat is enabled. Frag2Mat alone achieves $1.97-2.24\times$ speedup, and it still accelerates significantly with EarlyCull enabled. Therefore, Frag2Mat with G2L contributes the main speedup.
Table 5. Rendering FPS across settings, renderers, and datasets. To evaluate the effect of each component in TC-GS, we conducted an ablation study on EarlyCull and Frag2Mat with G2L.
EarlyCull Frag2Mat with G2L 3DGS SpeedySplat
truck drjohnson truck drjohnson
169.834 108.123 366.457 349.673
210.110 163.967 406.434 371.605
351.965 242.993 732.008 688.217
361.839 253.770 727.647 701.254
To clearly illustrate the impact of G2L on rendering quality and performance, we ran the following comparisons shown in Table 6 and the rendering images are shown in Figure 6. The ablation experiment shows that all Tensor Core variants boosted FPS. However, direct TF32 computation already introduced noticeable image-quality degradation, and FP16 alone dropped PSNR to just 8 dB. In contrast, combining FP16 Tensor Core optimization with Local Coordinates method preserved nearly same image quality while delivering the highest FPS. This demonstrates that our G2L can simultaneously maximize performance and maintain rendering fidelity.
Table 6. Ablation study on applying TC-GS on original 3DGS.
Frag2Mat G2L Precision drjohnson truck
$uparrow PSNR\uparrow FPS\uparrow$
@@ 5.3. Ablation Study
Figure 6. Ablation study on G2L when applying TC-GS on original 3DGS.
5.4. Discussion
5.4.1. Limitations